What is the surface area of the sidewalk if a city park is 400 'long & 300' wide. The sidewalks are 6' wide?
How do you find the ratio of the sidewalk to the area of the park?

Answer by rothauserc(4718)

surface area of park (SAP) = 400 * 300 = 120000 square feet
surface area of park with sidewalk (SAPW) = 412 * 312 = 128544
surface area of sidewalk = SAPW - SAP = 128544 - 120000 = 8544 square feet
ratio of sidewalk area to park area = 8544 / 120000 = 267 / 3750, approx 7/100